OpenStudy (anonymous):

log(x^3y^2)-2logx*(cubed root of y) - 3log (x/y)

OpenStudy (anonymous):

\[logx ^{3}y^{2}+2logx*\sqrt[3]{y}+3\log \frac{ x }{y }\]

OpenStudy (anonymous):

seems u want to simplify in order to solve, we will use the property of log \[\log(a^{m})= m* loga\] \[\log(a*b) = \log a + \log b\] \[\log(\frac{ a }{ b }) = loga-logb\]

OpenStudy (anonymous):

it says write the expression in one logarithm sorry

OpenStudy (anonymous):

in order to write in one logarithm , u have to use property of log i would suggest you to separate the x and y term first or in simple terms just do the simplification first because in the end it will be easy for you to write in single logarithm for example log(x^3*y^2) = log(x^3) + log(y^2) = 3logx + 2log(y) similarly do it for the other two terms ( using the appropriate property)

OpenStudy (anonymous):

how do I do it for \[-2logx\sqrt[3]{y}\] ?

OpenStudy (anonymous):

\[-2logx - 2\log(y^\frac{ 1 }{ 3 }) \]

OpenStudy (anonymous):

so the 3log x cancels and we have 2logy-3logy-2logx-2logy^1/3